OpenStudy (anonymous):

If the wavelength of a yellow light is 580nm what is its frequency?

OpenStudy (anonymous):

Recall that \[\nu = \frac{c}{\lambda}\]

OpenStudy (anonymous):

I get 5.169x10^-2

OpenStudy (anonymous):

I just wanted to check to see if I was working the problem correctly I don't entirely trust my calculator

OpenStudy (anonymous):

Your exponent is wrong. Up to orders of magnitude you had 10^8/10^-7 = 10^14

OpenStudy (anonymous):

thanks..... I also shouldn't trust my ability to convert apparently

OpenStudy (anonymous):

Make sure you recall nanometers are 10^-9 not 10^9 which may have caused that error

OpenStudy (anonymous):

I used \[580nm \times \frac{ 10^{-9}m }{ 1nm }\]